# WaterPark: A Robustness Assessment of Language Model Watermarking

### v1

@@ 5.2 Robustness – Observational Study
Next, we systematically assess the resilience of existing LLM watermarkers against representative watermark removal attacks listed in Table 3. Our main findings are summarized in Table 4, demonstrating the impact of a watermarker’s various design choices on its robustness to various attacks.
Design Choice Attack Category
Linguistic var. Lexical edit. Text-mix. Paraphras.
Index-dep.
Text-dep.
Context-free
Dist.-shift
Dist.-reweight
Dist.-transform
Score-based
Diff.-based
Model-based
Table 4: Impact of a watermarker’s design choices on its attack robustness ( positive; negative; neutral)
Figure 2: Resilience of LLM watermarkers against linguistic variation attacks.
@@ 5.2.1 Linguistic Variation Attack
This class of attacks perturb the linguistic features of the watermarked text $\tilde{T}$, without changing its semantics. Here, we consider three attacks: lowercasing, contracting, and expanding, which are concretely defined in §3.4. We measure the TPRs of different watermarkers against such attacks (with their FPRs fixed as 1%). Further, to illustrate the impact of the watermarker design on its attack resilience. We group the results along three design choices: context dependency, sampling strategy, and detection strategy, as shown in Figure 2.
@@
Remark 1 – Watermarkers that rely on explicit watermark signals tend to be robust against linguistic variation attacks.
@@ 5.2.2 Lexical Editing Attack
Figure 3: Resilience of LLM watermarkers against lexical editing attacks.
This class of attacks modifies individual words while maintaining the watermarked text’s semantics. Specifically, we consider four attacks in this category: synonymizing, misspelling, typoing, and swapping (see §3.4 for their concrete definitions). Similar to §5.2.1, we compare the TPRs of different watermarkers under these attacks (with their FPRs fixed as 1%). Figure 3 summarizes the results.
i) In terms of the context dependency design, text-dependent watermarkers (e.g., TRGL, UB, and GO) tend to be more vulnerable to lexical editing attacks, compared with index-dependent (e.g., RDF) and context-free (e.g., UG) watermarkers. Intuitively, as text-dependent watermarkers use previous tokens as the context for the next token in both the watermark generator and detector, the lexical editing thus causes a mismatch between the generator and detector. ii) Among the text-dependent watermarkers, it is observed that UB exhibits much higher vulnerability to such attacks, compared with the others. For example, its TPR drops near zero under the typoing and swapping attacks. This may be explained as follows. Recall that, to achieve unbiasedness, UB applies “hard” perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens tends to cause a more significant mismatch between the generator and detector, compared with other watermarkers that employ “soft” perturbation (e.g., distribution shift and transform). iii) The typoing attack is particularly effective, as it may cause significant tokenization errors.
Remark 2 – Text-dependent, (hard) distribution-reweight watermarkers are more vulnerable to lexical editing attacks.
@@ 5.2.3 Text-Mixing Attack
Figure 4: Resilience of LLM watermarkers against text-mixing attacks.
This class of attacks “dilutes” the watermark by mixing the watermarked text with non-watermarked text fragments. Here, to evaluate the resilience of different watermarkers against text-mixing attacks, we use the copy-pasting attack[27] as the concrete attack, which embeds the watermarked text into the context of non-watermarked, human-written text (generated under the same prompt). We use CP-$n$-$m$ to denote the attack in which the modified text $T^{\prime}$ consists of $n$ segments of watermarked texts, each of length $m$of $|T^{\prime}|$, and the rest as non-watermarked text. We consider 4 different settings to assess the impact of the lengths and positions of the watermarked text segments. The results are shown in Figure 4.
i) Overall, most watermarkers experience significant TPR drops, especially under CP-$1$-$10$ that only preserves 10% of the watermarked text. ii) Among all the watermarkers, GO and RDF show significantly higher attack resilience. This can be attributed to their sampling strategy: both employ distribution transform, which generates the next token deterministically conditional on a given random permutation. Thus, GO and RDF tend to have stronger per-token signals than the other watermarkers that sample the next token from a given pool (e.g., green list). This observation is consistent with that in §B.2. iii) Meanwhile, UB and UPV are the most vulnerable to the copy-pasting attack, with close to zero TPRs under CP-$1$-$10$ and CP-$1$-$25$. This can be explained as follows. The model-assisted detector of UPV determines the given text as watermarked based on its aggregated features (rather than per-token statistics), while the injected non-watermarked segments may greatly disrupt such features. Meanwhile, UB applies hard perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens causes a significant mismatch between the generator and detector.
Remark 3 – Distribution transform-based watermarkers show greater resilience to text-mixing attacks, while model-assisted, (hard) distribution-reweight watermarkers are more vulnerable.
Figure 5: Resilience of LLM watermarkers against paraphrasing attacks.
@@ 5.2.4 Paraphrasing Attack
This class of attacks employs an additional LLM (i.e., paraphraser) to re-write the given watermarked text $\tilde{T}$ (while preserving its semantics) to evade the detector. Here, we consider Dipper[28] as the paraphraser that rewrites $\tilde{T}$ in one shot. Specifically, we consider two settings of Dipper that respectively rewrite 20% and 40% of $\tilde{T}$. Further, we also consider the translating attack uses a translator model Seamless-m4t-v2-large[6] that first translates $\tilde{T}$ to French and then translates it back. Figure 5 compares the attack resilience of different watermarkers.
i) Both UPV and UB show higher vulnerability to the Dipper attack, compared with other watermarkers. For instance, the TPR of UPV drops to around zero under DP-$40$. This can be attributed to the same logic discussed in §5.2.3: the paraphrased text segments may greatly disrupt the aggregated textual features for UPV’s model-assisted detector, while the disruption to the previous tokens may cause a substantial mismatch between UB’s generator and detector, due to its rigid perturbation to the next-token distribution (e.g., rejecting half of the vocabulary). ii) Among the remaining watermarkers, RDF and UG are especially robust against the Dipper attack. This can be attributed to their index-dependent and context-free nature, which is less sensitive to the change of previous tokens than text-dependent watermarkers (e.g., TGRL, SIR, and GO). iii) Interestingly, RDF is more vulnerable to the translating attack than the Dipper attack. This can be explained by that RDF uses a detection strategy highly sensitive to the text length, while the translating attack can significantly shorten the text length.
Remark 4 – Index-dependent and context-free watermarkers tend to be more robust against paraphrasing attacks.
@@ 5.2.5 Fidelity Preservation
Recall that besides their attack effectiveness, another key metric for watermark removal attacks is whether they can preserve the quality of original texts. We thus compare the semantics of watermarked text $\tilde{T}$ and modified text $T^{\prime}$ using the metrics in §4. Figure 6 illustrates the quality preservation of different attacks on GO, with similar results on other watermarkers, with more results deferred to §D.3.
Figure 6: Quality preservation of different attacks on GO.
@@ 5.3 Robustness – Causal Analysis
Figure 7: Resilience of LLM watermarkers against multi attacks. a) Different in Context Dependency: TGRL (text-dependent) and UG (context-free) b) Different in Generation Strategy: TGRL (distribution-shift) and GO (distribution-transform) c) Different in Detection Method: UPV (Model-based) and UPV-key (Score-based)
@@ 5.3.1 Context Dependency
We select TGRL and UG to represent text-dependent and context-free designs respectively. TGRL uses the previous $k$ tokens as a randomness seed to divide the vocabulary into red/green lists for the current token, whereas UG uses fixed red/green lists in generating all the tokens. In general, UG performs similarly to TGRL. However, when subjected to the Dipper attack at varying intensity levels, UG consistently outperforms TGRL. For instance, as we increase the attack intensity (e.g. DP-60-20, DP-60-40), which involves rearranging the text, UG maintains a relatively stable TPR with only minor reductions. In contrast, TGRL’s TPR significantly drops, particularly at DP-60-20. Recall that the Dipper attack extensively paraphrases and reorders the text, leading to substantial changes in consecutive tokens. Due to its text-dependent design, TGRL struggles to maintain consistency in random seeds between detection and generation. In contrast, UG’s context-free design avoids this issue. This aligns with the results in UG[76], showing that when subjected to paraphrasing attacks, UG consistently outperforms TGRL.
@@ 6 Discussion
Next, we compare the attack resilience of watermarker-specific and generic detectors. For each watermarker, we apply the Dipper-$40$ attack and examine whether two detectors can effectively detect watermarked texts after the paraphrasing attack. Figure 8 depicts the confusion matrices of both detectors.
@@
Remark 5 – Employing a generic detector alongside a watermark-specific detector improves the security of vulnerable watermarkers.
@@ 6.2 Advanced Attack
In addition to the previously mentioned simple attacks, we investigate the effectiveness of more advanced attacks.
@@ 6.2.1 Varying Attack Intensity
One straightforward way to improve an attack’s effectiveness is to increase its intensity, potentially at the cost of other metrics (e.g., text quality). Here, we consider the Dipper attack[28] under varying intensity settings, denoted as DP-$l$-$o$, where lexical change ($l$) indicates that $l$of the given text is paraphrased and order change ($o$) indicates that $o$of the text is re-ordered. We compare the watermarkers’ resilience under varying attack intensity, as shown in Figure 9.
i) As expected, most watermarkers observe TPR drops as the attack intensity increases. ii) Among all the watermarkers, UG demonstrates the most consistent resilience under varying attack intensity. This can be attributed to its context-free design: the same perturbation is applied to the next-token distribution across all the tokens, which is thus immune to the change of previous tokens. iii) Compared with text-dependent watermarkers (e.g., TGRL and GO), an index-dependent watermarker (e.g., RDF) shows stronger resilience, especially under high attack intensity (e.g., DP-$60$-$20$), due to its weaker dependency on previous tokens. iv) UPV’s performance is somewhat inconsistent; it struggles with low-intensity attacks (e.g., DP-20) but shows resilience to high-strength ones (e.g., DP-60). This inconsistency can be attributed to UPV’s model-assisted detector and the inherent instability of its neural network, as confirmed by repeated experiments.
Remark 6 – A watermarker’s resilience to increasingly intensive attacks varies with its context dependency, roughly, context-free $>$ index-dependent $>$ text-dependent.
@@ 6.2.2 Combining Simple Attacks
We first explore whether combining two attacks improves the attack’s effectiveness. Here, we only consider combining two simple attacks from the “weak” linguistic variation and lexical editing attacks. Figure 10 illustrates the effectiveness of such combined attacks.
@@
Remark 7 – Evaluating watermarkers’ robustness needs to consider not only individual attacks but also their combinations.
@@ 6.2.3 Adaptive Attack
Given the availability of generic detectors, it is possible for the adversary to exploit such detectors to adapt their attacks. We consider an attack scenario as shown in Figure 11: the adversary performs a gradient-based adversarial attack[18] that iteratively modifies the watermarked text to evade the generic detector, and then forwards the modified text to the target, watermark-specific detector.
Figure 12: Detection of watermarked texts by watermarker-specific and generic detectors on gradient-based attacked samples (‘1’ or ‘0’ indicate that the detector detects the given watermarked text as watermarked or non-watermarked).
We evaluate this attack’s effectiveness using 500 watermarked texts from the C4 dataset. We use GBDA[18] as the adversarial attack and limit the number of perturbations to 100. The results are summarized in Figure 12. i) Leveraging the surrogate detector significantly improves the attack effectiveness: with the BERTScore between $\tilde{T}$ and $\tilde{T}^{\prime}$ is about 0.76, slightly lower than the Dipper attack, the detection rates of most watermarkers drop below 10%. ii) Although some samples do not evade the generic detector, they evade the specific detector successfully (e.g., TGRL, UG, and UB). iii) UG and RDF exhibit greater robustness than the other watermarkers. Specifically, for RDF, 42% of the samples evade the generic detector but are still detected by the RDF-specific detector. This superior robustness is likely due to their weaker context dependencies.
Remark 8 – Evaluating a watermarker’s robustness needs to consider attacks that leverage surrogate detectors.
@@ 6.2.4 Leveraging Expert LLMs
We now explore the question of “what if the adversary has access to highly capable LLMs?” Specifically, we implement another paraphrasing attack that employs ChatGPT-3.5-Turbo as the paraphraser. For each watermarker, we randomly sample 100 watermarked texts that are successfully detected, and query the ChatGPT API with the following prompt:
@@
and then forward the paraphrased text to the detector. To further evaluate the impact of the paraphrasing strength, we also measure the attack effectiveness under multiple rounds of paraphrasing as suggested in[72]. Notably, the attack described in [72] employs the T5 model to paraphrase the entire text up to 300 times. In contrast, our study limits paraphrasing to a maximum of 5 iterations. The difference between these threat models is due to the distinct objectives of the two studies: while [72] aims to understand the lower bound of watermark robustness, we focus on evaluating watermark robustness against practical, resource-limited adversaries. The results are summarized in Figure 13.
@@
We have the following findings. i) The detection rates of all the watermarkers drop below 0.3 after one round of GPT paraphrasing, in which SIR and UB become nearly undetectable. This indicates that using highly capable LLMs to paraphrase watermarked texts is a dominant attack that effectively nullifies most existing watermarkers. However, as specified in our threat model (§2.3), access to highly capable LLMs may fall outside the scope of our robustness assessment, as their availability negates the need for watermark removal attacks. ii) For watermarkers that survive the first round of paraphrasing (e.g., UG and GO), their detection rates quickly drop below 0.15 as the adversary applies multiple rounds of paraphrasing. This observation corroborates the findings in[72].
Remark 9 – Paraphrasing using highly capable LLMs serves as a dominant attack to test watermarkers’ robustness lower-bounds.
6.3 Evaluation Guidelines
Robustness. It is often critical to properly set the decision threshold for a watermarker (and also the attacks) to fully assess its robustness[41], which unfortunately is often missing in the original papers. To overcome this issue, we recommend i) measuring the overall effectiveness (TPR) in terms of ROC (across different threshold settings), ii) measuring the TPR under a fixed FPR (e.g., 0.01), and iii) considering varying attack intensity.
Fidelity. It is notoriously challenging to meaningfully measure the quality of text data[50]. We recommend employing a variety of metrics (e.g., BERTScore, P-SP, MAUVE) to comprehensively measure the quality retention of watermarkers as well as attacks. In addition, one may also leverage external tools (e.g., more advanced LLMs such as GPT-4) or human evaluation to provide a more accurate assessment if feasible.
7 Related Work
Next, we survey the relevant literature in the following categories: i) detection of LLM-generated texts, ii) LLM watermarking, iii) attacks on LLM watermarking, and iv) evaluation of LLM watermarkers.
Detection of LLM-generated texts. The advances in LLMs also give rise to their possible misuses[64, 63, 62, 61, 77, 16, 24, 9, 71, 33, 38, 23]. There is thus a pressing need for the capability of distinguishing LLM- and human-generated texts. Initial work attempts to either train classifiers using LLM- and human-generated texts[44] or to leverage intrinsic characteristics of LLM-generated texts (e.g., perplexity and variability in length, complexity, and information density)[42]. Yet, with LLMs becoming increasingly capable, the difference between LLM- and human-generated texts is narrowing, making such approaches less effective.
LLM watermarking In response, LLM watermarking emerges as a promising alternative, which instruments the LLM generative process with statistical signals that can be subsequently detected. The existing LLM watermarking techniques can be categorized based on the stages in which they are applied[37]: i) training-time watermarking[40, 57, 56, 55, 67], ii) watermarking during logit-generation[26, 76, 20, 35, 36, 12, 51, 14, 60, 69, 31, 17], iii) watermarking during token-sampling [29, 1, 19], and iv) post-generation watermarking[73, 70, 68, 45, 53, 2]. This study mainly focuses on training-free, pre-generation watermarking, which applies to any given LLMs and provides flexible control over multiple criteria (e.g., quality, effectiveness, and robustness).
Attacks on LLM watermarking. One critical property of an LLM watermarker is its robustness against potential attacks. A variety of attacks can be applied to LLM watermarking, ranging from removing the embedded watermark to uncovering the green/red lists. For instance, Dipper[28] is a widely used paraphrasing attack to evaluate the robustness of LLM watermarkers [26, 76] against watermark removal attacks; the watermark stealing attack[25] is proposed to identify green-list tokens and to replace them with red-list tokens, targeting TGRL[26] and UG[76] to further launch spoofing or removal attacks. This study primarily focuses on watermark removal attacks as they can target any LLM watermarkers and have profound implications in practice (e.g., disinformation, academic cheating, and automated phishing).
Evaluation of LLM watermarkers. As LLM watermarkers become prevalent, recent work attempts to benchmark the performance of various watermarkers. However, current studies either focus solely on the effectiveness of watermarking or have limited assessments of robustness. For instance, WaterBench[59] compares the effectiveness of TGRL[26] and UG[76] under varying hyper-parameter settings (e.g., prompt length); LLM-judger[54] uses GPT-3.5-Turbo as a judge to evaluate the effectiveness of RDF[29] and TGRL[26] and employs a binary classifier based on MLP to distinguish between watermarked and non-watermarked texts; MarkMyWords[49] compares the effectiveness of four watermarkers including TGRL[26], GO[1], RDF[29], and UW[5] and only evaluates the robustness of TGRL against watermark removal attacks.

### v2

@@ 4.2 Robustness – Observational Study
We systematically assess the resilience of existing LLM watermarkers against representative watermark removal attacks. All the results are shown in Table  3.
Water marker CLEAN Linguistic variation Lexical editing Text-mixing Paraphrasing
Contra Expan LowCase Swap Typo Syno Missp CP1-10 CP3-10 CP1-25 CP3-25 DP-20 DP-40 Trans
TGRL 0.992 0.994 0.994 0.984 0.874 0.788 0.996 0.992 0.034 0.176 0.126 0.836 0.952 0.858 0.652
UG 0.964 0.963 0.965 0.879 0.956 0.884 0.965 0.956 0.131 0.219 0.192 0.534 0.940 0.916 0.658
UPV 0.776 0.772 0.814 0.685 0.667 0.355 0.764 0.695 0.050 0.073 0.242 0.450 0.296 0.057 0.006
RDF 0.996 0.993 0.993 0.983 0.951 0.961 0.990 0.993 0.094 0.396 0.430 0.974 0.988 0.962 0.154
UB 1.000 1.000 1.000 0.946 0.030 0.088 0.992 0.996 0.002 0.009 0.048 0.410 0.461 0.149 0.306
SIR 0.954 0.954 0.954 0.886 0.932 0.566 0.936 0.904 0.138 0.210 0.178 0.466 0.902 0.814 0.818
GO 0.998 0.998 0.998 0.988 0.920 0.852 0.996 1.000 0.181 0.589 0.620 0.992 0.971 0.847 0.928
Table 3: Attack resilience of LLM watermarkers. The intensity of red shading indicates higher values, while the intensity of blue shading indicates lower values, with 0.5 serving as the threshold between the two color gradients.
@@ 4.2.1 Linguistic Variation Attack
This attack perturbs the linguistic features of the watermarked text, without changing its semantics.
@@ 4.2.2 Lexical Editing Attack
This attack modifies individual words while maintaining the watermarked text’s semantics.
i) TRGL, UB, SIR and GO tend to be more vulnerable to lexical editing attacks, compared with RDF and UG watermarkers. Intuitively, as text-dependent watermarkers (e.g., TRGL, UB, and GO) use previous tokens as the context for the next token in both the watermark generator and detector, the lexical editing thus causes a mismatch between the generator and detector. ii) UB exhibits much higher vulnerability to such attacks, compared with the others. For example, its TPR drops near zero under the typoing and swapping attacks. This may be explained as follows. Recall that, to achieve unbiasedness, UB applies “hard” perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens tends to cause a more significant mismatch between the generator and detector, compared with other watermarkers that employ “soft” perturbation (e.g., distribution shift and transform). iii) The SIR shows notably inferior performance, especially when subjected to typoing attacks. While it employs a neural network to predict token-specific logit perturbations that are designed to be unbiased (no preference for particular tokens) and balanced (with perturbations summing to zero), its results are suboptimal, suggesting that it is fundamentally challenging to make accurate predictions under such perturbations. iv) The typoing attack is particularly effective, as it may cause significant tokenization errors.
@@ 4.2.3 Text-Mixing Attack
This class of attacks “dilutes” the watermark by mixing the watermarked text with non-watermarked text fragments. Here, to evaluate the resilience of different watermarkers against text-mixing attacks, we use the copy-pasting attack (Kirchenbauer et al., 2023b) as the concrete attack, which embeds the watermarked text into the context of non-watermarked, human-written text (generated under the same prompt). We use CP-$n$-$m$ to denote the attack in which the modified text $T^{\prime}$ consists of $n$ segments of watermarked texts, each of length $m$of $|T^{\prime}|$, and the rest as non-watermarked text.
i) Overall, most watermarkers experience significant TPR drops, especially under CP-$1$-$10$ that only preserves 10% of the watermarked text. ii) Among all the watermarkers, GO and RDF show significantly higher attack resilience. This can be attributed to their sampling strategy: both employ distribution transform, which generates the next token deterministically conditional on a given random permutation. Thus, GO and RDF tend to have stronger per-token signals than the other watermarkers that sample the next token from a given pool (e.g., green list). This observation is consistent with that in §E.2. iii) Meanwhile, UB and UPV are the most vulnerable to the copy-pasting attack, with close to zero TPRs under CP-$1$-$10$ and CP-$1$-$25$. This can be explained as follows. The model-assisted detector of UPV determines the given text as watermarked based on its aggregated features (rather than per-token statistics), while the injected non-watermarked segments may greatly disrupt such features. Meanwhile, UB applies hard perturbation on the next-token distribution (e.g., by rejecting half of the vocabulary); thus the disruption to the previous tokens causes a significant mismatch between the generator and detector.
@@ 4.2.4 Paraphrasing Attack
This class of attacks employs an additional LLM (i.e., paraphraser) to re-write the given watermarked text $\tilde{T}$ (while preserving its semantics) to evade the detector. Here, we consider Dipper (Krishna et al., 2023) as the paraphraser that rewrites $\tilde{T}$ in one shot. Further, we also consider the translating attack uses a translator model Seamless-m4t-v2-large (Communication et al., 2023) that first translates $\tilde{T}$ to French and then translates it back.
i) UPV and UB exhibit higher vulnerability to the Dipper attack, compared to other watermarkers. UPV’s TPR drops to around zero under DP-$40$, which aligns with our analysis in §4.2.3: the paraphrased text segments may greatly disrupt the aggregated textual features for UPV’s model-assisted detector, while the disruption to the previous tokens may cause a substantial mismatch between UB’s generator and detector, due to its rigid perturbation to the next-token distribution (e.g., rejecting half of the vocabulary). ii) In contrast, RDF and UG are especially robust against the Dipper attack. This can be attributed to their index-dependent and context-free designs, which are less sensitive to the change of previous tokens than text-dependent watermarkers (e.g., TGRL, SIR, and GO). iii) Interestingly, RDF is more vulnerable to the translating attack than the Dipper attack. This susceptibility arises from RDF’s detection mechanism, which is highly sensitive to text length variations, a weakness readily exploited by the translating attack’s tendency to produce shorter output text.
@@ 4.2.5 Fidelity Preservation
Recall that besides their attack effectiveness, another key metric for watermark removal attacks is whether they can preserve the quality of original texts. We thus compare the semantics of watermarked text $\tilde{T}$ and modified text $T^{\prime}$ using the metrics in §3.2. Figure 2 illustrates the quality preservation of different attacks on GO, with similar results on other watermarkers, with more results in §G.3.
Figure 2: Quality preservation of different attacks.
@@ 4.3 Robustness – Causal Analysis
Figure 3: Watermarker robustness to multi-attacks. a) Context dependency: TGRL (text-dependent) and UG (context-free); b) Generation strategy: TGRL (distribution-shift) and GO (distribution-transform); c) Detection method: UPV (Model-based) and UPV-key (Score-based).
@@ 4.3.1 Context Dependency
We select TGRL and UG to represent text-dependent and context-free designs respectively. TGRL uses the previous $k$ tokens as a randomness seed to divide the vocabulary into red/green lists for the current token, whereas UG uses fixed red/green lists in generating all the tokens. In general, UG performs similarly to TGRL. However, when subjected to the Dipper attack at varying intensity levels, UG consistently outperforms TGRL. For instance, as we increase the attack intensity (e.g. DP-60-20, DP-60-40), which involves rearranging the text, UG maintains a relatively stable TPR with only minor reductions. In contrast, TGRL’s TPR significantly drops, particularly at DP-60-20. Recall that the Dipper attack extensively paraphrases and reorders the text, leading to substantial changes in consecutive tokens. Due to its text-dependent design, TGRL struggles to maintain consistency in random seeds between detection and generation. In contrast, UG’s context-free design avoids this issue. This aligns with the results in UG (Zhao et al., 2023), showing that when subjected to paraphrasing attacks, UG consistently outperforms TGRL.
@@ 5 Discussion
We compare the attack resilience of watermarker-specific and generic detectors. For each watermarker, we apply the Dipper-$40$ attack and examine whether two detectors can effectively detect watermarked texts after the paraphrasing attack. Figure 4 depicts the confusion matrices of both detectors.
@@ 5.2 Advanced Attack
In addition to the previous simple attacks, we investigate the effectiveness of more advanced attacks.
@@ 5.2.1 Varying Attack Intensity
One straightforward way to improve an attack’s effectiveness is to increase its intensity, potentially at the cost of other metrics (e.g., text quality). Here, we consider the Dipper attack (Krishna et al., 2023) under varying intensity settings, denoted as DP-$l$-$o$, where lexical change ($l$) indicates that $l$of the given text is paraphrased and order change ($o$) indicates that $o$of the text is re-ordered. We compare the watermarkers’ resilience under varying attack intensity, as shown in Figure 5.
i) As expected, most watermarkers observe TPR drops as the attack intensity increases. ii) Among these, UG demonstrates the most consistent resilience under varying attack intensity. This can be attributed to its context-free design: the same perturbation is applied to the next-token distribution across all the tokens, which is thus immune to the change of previous tokens. iii) Compared with text-dependent watermarkers (e.g., TGRL and GO), an index-dependent watermarker (e.g., RDF) shows stronger resilience, especially under high attack intensity (e.g., DP-$60$-$20$), due to its weaker dependency on previous tokens. iv) UPV’s performance is inconsistent; it struggles with low-intensity attacks (e.g., DP-20) but shows resilience to high-strength ones (e.g., DP-60). This inconsistency can be attributed to UPV’s model-assisted detector and the inherent instability of its neural network, as confirmed by repeated experiments.
@@ 5.2.2 Combining Simple Attacks
We first explore whether combining two attacks improves the attack’s effectiveness. Here, considering the most feasible combinations, we only combine two simple attacks from the “weak” linguistic variation and lexical editing attacks. Figure 6 illustrates the effectiveness of such combined attacks.
@@ 5.2.3 Adaptive Attack
Given the availability of generic detectors, it is possible for the adversary to exploit such detectors to adapt their attacks. We consider a scenario as shown in Figure 7: the adversary performs a gradient-based attack (Guo et al., 2021) that iteratively modifies the watermarked text to evade the generic detector, and then forwards the modified text to the target, watermark-specific detector.
Figure 8: Watermark detection by watermarker-specific and generic detectors on gradient-based attacked samples (‘1’ or ‘0’ indicate that the detector detects the given watermarked text as watermarked or non-watermarked).
We evaluate this attack’s effectiveness using 500 watermarked texts from the C4 dataset. We use GBDA (Guo et al., 2021) as the adversarial attack and limit the steps of perturbations to 100. Compared to Dipper, GBDA makes more substantial textual modifications. The results are summarized in Figure 8. i) Leveraging the surrogate detector significantly improves the attack effectiveness: with the BERTScore and BLEU-1 between $\tilde{T}$ and $\tilde{T}^{\prime}$ is about 0.764 and 0.188, respectively, slightly lower than the Dipper attack, the detection rates of most watermarkers drop below 10%. ii) Although some samples do not evade the generic detector, they evade the specific detector successfully (e.g., TGRL, UG, and UB). iii) UG and RDF exhibit greater robustness than the other watermarkers. Specifically, for RDF, 42% of the samples evade the generic detector but are still detected by the RDF-specific detector. This superior robustness is likely due to their weaker context dependencies.
@@ 5.2.4 Leveraging Expert LLMs
We have the following findings. i) The detection rates of all the watermarkers drop below 0.3 after one round of GPT paraphrasing, indicating that using highly capable LLMs to paraphrase watermarked texts is a dominant attack that effectively nullifies most watermarkers. However, as specified in our threat model (§3.1), access to highly capable LLMs may fall outside the scope of our robustness assessment, as their availability negates the need for watermark removal attacks. ii) For watermarkers that survive the first round of paraphrasing (e.g., UG and GO), their detection rates quickly drop below 0.15 as the adversary applies multiple rounds of paraphrasing. This observation corroborates the findings in (Zhang et al., 2023).